[Halla_running] [New Logentry] Hall A Status update -- Thurs AM

brads at jlab.org brads at jlab.org
Thu Mar 19 09:35:02 EDT 2015


Logentry Text:
--
<p>The Hall will remain in beam permit until roughly noon today. MCC will be trying to deliver beam to us to study/address the BLM trips at 1C12. The first thing to do when beam gets to our dump is to run the Harps and see if they are fixed, then proceed to a bull's eye scan.</p><p>Beam will be going off at noon so Accel can fix the east arc heat exchanger -- we expect to be down for 6--8 hours. (https://logbooks.jlab.org/entry/3327870)</p><p>We will go to <strong>controlled access</strong> when beam shuts off at noon. Check in with the Shift Leader before heading down to the Hall and <span style="text-decoration: underline;">please make a log entry</span> to describe what was done.</p><ul><li>The Techs have a few things they would like to look at.</li><li>Alignment wants t o pop in and take a look at the target chamber (they're trying to understand the source of the target misalignments identified earlier).</li></ul><p>I&amp;C will also be looking at the 1C18 harps (upstream
  of the compton in the Hall A arc).</p>

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3328064
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://mailman.jlab.org/pipermail/halla_running/attachments/20150319/de3a740e/attachment.html 


More information about the Halla_running mailing list